OpenStudy (anonymous):

write in vertex form: -4x^2 + 2x + 1 can someone remind me how to do this

OpenStudy (campbell_st):

y = -4(x^2 - 1/2x + 1/16) + 5/4 y = -4(x - 1/4)^2 + 5/4